OpenStudy (anonymous):

35 ÷ 5 + 0.32 × 4 - 1.12 35 ÷ 5 + 0.3(2 to the second power) × 4 - 1.12

OpenStudy (anonymous):

BEDMAS

OpenStudy (anonymous):

u mean pemdas?

OpenStudy (anonymous):

samething

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

yes they are

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

parentheses are also known as brackets, therefore one can say PEDMAS or BEDMAS. they are equal

OpenStudy (anonymous):

always if u have question first with parentheses if u dont have parentheses start with powres \[2^2\] next divide or Product .. after that sum or Difference

OpenStudy (anonymous):

\[35 ÷ 5+0.32 × 2^2 -1.12 = 35 ÷ 5+0.32 × 4 \]\[35 ÷ 5 =7\]\[0.32*4=1.28\] so we have .. \[7+1.28-1.12\]\[7+1.28=8.28\]\[8.28-1.12=7.16\]

OpenStudy (anonymous):

i hope i can help u .. ;D